ROTATIONAL AND RENNER-TELLER ANALYSES OF THE GROUND AND EXCITED STATE LEVELS OF THE JET-COOLED CS_{2}^{+} ION

\underline{DENNIS J. CLOUTHIER} and SHENG-GUI HE, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The Ã^{2}Π_{u} −X̃^{2}Π_{g} spectrum of the jet-cooled carbon disulfide cation (CS_{2}^{+}) has been studied by laser-induced fluorescence and single vibronic level emission spectroscopic techniques. The ions were produced in a pulsed discharge jet using a precursor mixture of CS_{2} vapor in argon. Analysis of the high-resolution spectrum of the ^{2}Π_{3/2} vibronic component of the 0-0 band has provided reliable ground and excited state molecular structures. A total of 33 LIF bands have been observed including a variety of hot bands from the μ^{2}Σ and ^{2}Δ_{5/2} vibronic components of the first bending level in the ground state. In emission, 34 ground state levels were observed, including all four components of v_{2} = 1. The ground and excited state vibronic energy levels were fitted using a Renner-Teller model that included spin-orbit, vibrational anharmonicity, and Fermi resonance interactions. The spin-orbit splittings are much larger in the ground that in the excited state, although the Renner-Teller constants ϵ are comparable. The ν_{1} − 2ν_{2} Fermi resonance interaction is very strong in both states and causes a nearly 50:50 mixing of the basis functions in each case.